VIBRATIONAL PREDISSOCIATION DYNAMICS OF HF-COMPLEXES AT v_{HF} =3

\underline{CHENG-CHI CHUANG}, SUSY N. TSANG, KELLY HIGGINS and WILLIAM KLEMPERER, Department of Chemistry, Harvard University, Cambridge, MA 02138; HUAN-CHENG CHANG, Institute of Atomic and Molecular Sciences, Academica Sinica, P.O.Box 23-166, Taipei, Taiwan 10764, Republic of China.

We report the lifetimes and rotational distributions of HF products from the vibrational predissociation (VP) of the ArHF and Ar_{3}HF at v_{HF} =3 by dispersing laser induced fluorescence. The product state distribution of these two complexes has shown that the excess energy resulting from VP (Δv = −1) is partioned primarily into HF(v = 2) product rotation. The distribution is extremely sharply peaked near the highest energetically accessible rotational channel (J′=13 for ArHF(3000) and J′=12 for Ar_{3}HF). The lifetimes of ArHF(3000) and (3110) states are found to be 4.8×10^{−6} sec and 8.9×10^{−6} sec respectively. Much shorter vibrational predissociation lifetime of 3 nsec is determined for Ar_{3}HF at v = 3, revealed from the Lorentzian component of 60 MHz of the Voigt profile. The correlation of VP lifetime and red-shift for HF-complexes at v = 3 will be compared to that at v = 1. At the fundamental valence excitation (v_{HF} = 1 ← 0), the VP rate varies as the square of the frequency red-shift.^{a} At the present second overtone excitation (v_{HF} = 3 ← 0), based upon (HF)_{2}, N_{2}HF and Ar_{3}HF results, the VP rate varies more dramatically as the cube of the frequency red-shift. This suggests that a relatively complex model is required to fully account for the enhancement of VP rates observed for HF-complexes at v = 3.

aR. E. Miller, Science 240, 447 (1988)